OpenStudy (anonymous):

Find the ratio in simplest form. 2/3 to 3/2 A. 9/4 B. 1 C. 4/9

OpenStudy (anonymous):

(2/3)/(3/2) and solve...

OpenStudy (anonymous):

divide?

OpenStudy (anonymous):

Yes, ration basically means divide.

OpenStudy (anonymous):

*ratio

OpenStudy (anonymous):

well no instead you can now multiply the top and bottoms of this complex fraction to reduce it to a simpler fraction

OpenStudy (anonymous):

for example (1/2)/(1/2) multiply tom and bottom fraction by 2 ((1/2)*2)/((1/2)*2) =2/2 =1 see?

OpenStudy (anonymous):

well the both answer 6?

OpenStudy (anonymous):

no 6 is not the answer you need to simplify the complex fraction (2/3)/(3/2) multiple top and bottom by3 multiply top and bottom fraction by 2 vwala simplified

OpenStudy (anonymous):

its 4/9? i guessed cause i dont have a calculater that will do fractions.

OpenStudy (anonymous):

YAY!!! (=

OpenStudy (anonymous):

good job!

OpenStudy (anonymous):

thanks alot cuz im retarded :) lol

OpenStudy (anonymous):

ha ha, no just make sure you remember that you can simplify fractions that way... hopefully you learned something...

OpenStudy (anonymous):

yup:)